Point charges of -6 mue C, and +6 mue C are locate
Answer #1

E.F. at x=3 cm due to -6\muC , E1= -6x10-6x9x109/ (4x10-2)2 N/C

E.F. at x=3 cm due to 5\muC , E2= 5x10-6x9x109/ (3x10-2)2 N/C

E.F. at x=3 cm due to 6\muC , E3= 6x10-6x9x109/ (2x10-2)2 N/C

Total E.F. at x=3 cm, E = E1+E2+E3 = 15.12 x 107 N/C

E.F. at x=15 cm due to -6\muC , E4= -6x10-6x9x109/ (16x10-2)2 N/C

E.F. at x=15 cm due to 5\muC , E5= 5x10-6x9x109/ (15x10-2)2 N/C

E.F. at x=15 cm due to 6\muC , E6= 6x10-6x9x109/ (14x10-2)2 N/C

Total E.F. at x=15 cm, E = E4+E5+E6 = 0.27 x 107 N/C

Say at x cm the E.F. is zero.

so we can write

9x109x 10-6 x104[-6/(x+1)2+5/x2+6/(x-1)2]=0

Solving this we get x=0.69 cm